a plane which can fly 520mph in still air flies for 3 hours against a wind and for 2 hours with the same wind. The total distance it covers is 2565 miles. find rate of the wind.
Answer by Mathtut(3670)

lets call s the rate of the plane which is 520mph d=2565
now with the formula distance(d)=rate*time or d=rt lets break this trip down into 2 segments. We will call the speed of the wind w.
d=(s-w)3+(s+w)2---->2565=(520-w)3+(520+w)2--->2565=1560-3w+1040+2w
w=35 after combining like terms and solving for w.
so the rate of the wind 35mph